last modified… 9
- 9 NOV
(neff): There have been some minor modifications
to the checklist as a result of today's training session. Added a few items to the
Also note that the TCS program and TheSky
start up when you log in to the observer account. If the dome is not pointing north
and the telescope at the zenith once TCS starts, you can move them to
their "home" positions with the handpaddle
and then restart TCS.
- 8 NOV
Students in ASTR377 will be doing various projects in the dome
during November. There might
be instruments around, the computer might be logged into a different
account, and the telescope might not be configured for the eyepiece. If you wish to use the telescope
this month, it would be a very good idea to check with Neff in advance.
- 14 OCT
Use the new CHECKLIST.
Supplemental guides for using the imager and spectrometer are being
prepared. There are 3 generic
accounts: use "OBSERVER" if you just plan to use the eyepiece;
use "IMAGER" if you are using the Apogee AltaU6 ccd camera; use "SPECTROMETER" if you are
using the SBIG Self-Guiding Spectrograph.
- 12 OCT
12 (neff): Dome shutter power switch has been
- 10 OCT
12 (neff): Dome shutter power switch is flakey
to inoperable. Don't open
unless you know you can close the shutter manually before any potential
-------------------------------------------- (archives; most is not relevant any
- 7 NOV
The observer account still doesn’t work, and the start up/shutdown
scripts will bomb because the dome is not run from the computer right
now. So log in as CCD and
just run the sky then pull down “Telescope/Link/Establish in the Sky after
you’ve got TCS set up.
- 1 NOV
The PC serial line has been removed. The dome control works fine manually. The OBSERVER account is not
properly configured, and the orchestrate script will bomb. I plan to fix this before the
public night Friday. The CCD
account is properly configured.
I will be taking the CCD off for public night.
- 31 OCT
We are having problems with the dome azimuth controller. I will be fixing it Wednesday and
Thursday. In the mean time,
contact me if you need to use the telescope.
- 11 APR
I put the CCD back on the filterwheel/pickoff
mirror. We don’t have enough
slack in the USB cable to safely use it with the 16”, and we don’t have a
long serial cable for the filter wheel yet. It is possible to mount it on the scope now, but let’s
wait until I can get some longer cables. In the meantime, Charles Taylor is doing a lot of tests
with the CCD in the control room, so make sure that nothing is running
before you try to observe.
- 11 APR
The filter wheel and dome stopped working because the port
definitions in the Keyspan Serial->USB box
got changed. I don’t know
what causes that, but it’s happened before. I reset everything to the way they are supposed to be
(dome on port1/COM1; tcs computer on port3/COM6;
filterwheel on port4/COM3). The long serial cable definitely
is hosed, but it will be replaced any day now. Anyway, everything works again, but don’t try to use
the B or V filter. In fact,
leave it on position 1 (empty hole) unless you have a reason to do
otherwise. Better still,
don’t connect to it at all (select “none” rather than “DFM filter wheel”
in CCDSoft setup menu. Please let me know if any problems develop.
- 7 APR
07 (neff): The filter wheel now
works again. I will get a new
serial cable, then we will clean the filters and
re-attach the CCD so it can be mounted on the telescope. This should be ready by early next
week. The bad news is that
the B and V filters are totally ruined. They will be replaced over the summer. Students are doing 370 projects
right now, so be very careful not to disturb their work for the next few
- 20 MAR
The new CCD camera works fine. Now we are having trouble getting the filter wheel to
respond. We have also
isolated, but not yet fixed, the problem with the dome azimuth
controller. Progress is being
made. Telescope is usable,
without the CCD/filter wheel.
If you need to mount the camera on the telescope, let me know.
- 2 MAR
Collimation looks very good!
Telescope is ready for use with eyepiece. RTL and I tried to plug the new CCD in, but it will
take some TLC to make it work.
After spring break we will get back to work on it. Have fun with the Lunar Eclipse!
- 1 MAR
Telescope is back in service.
Please monitor the collimation and let me know if it needs to be
adjusted. The CCD has been
replaced and is not yet ready to return to service. Filters will be cleaned before we
put the filterwheel back on the telescope. Happy observing!
- 13 DEC
Observatory is DOWN until further notice. The primary mirror has been removed. It will be sent off for recoating
and reinstalled in about a month.
THE CCD camera and dome azimuth controller will be fixed while we
are waiting for the mirror.
Hopefully everything will be back in place for next semester. Have a good holiday break.
- 29 SEP
06 (neff): The dome controller
is not working, but it still needs to be plugged in while the telescope is
working (The Sky needs feedback from Automadome). Please turn the dome controller
off, disconnect the PC serial line, and shut off the UPS to the dome
control when you are finished.
- 26 SEP
The CCD has been removed from the telescope. It is still available for any
student projects, but we need to dry it out before it is reinstalled. Please let me know if you plan to
use it; in the mean time we’ll leave the eyepiece on there
for labs and public viewing.
- 20 JUN
Filter wheel is back on-line.
It was set to com port 2. I reset it to com
port 8, and everything works.
The mapping on the keyspan is such that
control (Automadome); #2=COM6èTCS
Computer (The Sky); #3=COM7 (empty); #4=COM8èDFM
Filter Wheel (controlled by CCDSoft). If things get changed arbitrarily
again, make sure the definitions in the software set-up menus match these
mappings. Everything is
working except the dome azimuth controller (see 10 Feb) message.
- 10 Feb
06 (neff & true): Telescope is back on-line. There are still problems, so read on…. We installed the roller in the
dome track and greased the tracks and gears. We also re-installed the dome azimuth controller. The dome now works very well, and
the azimuth control works fine manually (using the right and left buttons
in the dome). However, the
dome controller has lost all its memory, so it is not working with Automodome.
We are looking into this, but the scope is usable now (you just
have to run upstairs and tweak the dome pointing after each move). We also took a hammer to the filter
wheel, and it seems to behave for us. Actually, we fiddled with the Keyspan
control software and the DFM filter wheel setup in CCDSoft. Anyway, it works now.
- 3 MAR
05 (neff & true): We fixed the serial cable for the filter wheel. EVERYTHING WORKS NOW!! The next CCD user needs to route
the cables up to the dome before mounting the camera. We still need a more permanent
- 23 FEB
The filter wheel works fine with a short cable. I’ll go get us a good serial cable
and hook it up so that everything will work right with the ccd. Any CCD users should speak with me
about telescope balance before and after using the camera. We are working on a more permanent
solution. In the meantime,
everything is working ok.
- 17 NOV
Telescope should work now.
COM1 is the dome, COM6
is the TCS. CCD Users
will have to figure out how to configure the filter wheel serial
port. All serial ports now go
into the Keyspan serial->USB adapter. The filter wheel runs only at 1200
Baud. The others run at 9600.
- 11 NOV
Telescope is down until further notice. The dome controller has been sent back to the factory
for a re-wiring. As soon as
it gets back, I’ll install it and we’ll be back in business. Also in the meantime, we’ll get
some more serial ports.
Expect down time of about 2 weeks.
- 9 NOV
We seem to have lost yet another serial port. I robbed the serial-to-usb line from the filter wheel (com2) and plugged the
TCS computer into com2. It is
now no longer possible to run the telescope using TheSky
simultaneous with using the camera and filter wheel. If you are a CCD observer and want
to use the filter wheel, simply disconnect the TCS line from com2 and plug
the filter wheel line into it (you don’t have to change any
settings). Run the telescope
using TCS, not TheSky. Run the CCD with CCDSoft. We’ll get some new ports asap.
- 5 NOV
The telescope works fine, but the computer control of the dome
rotation does not. The dome
can be rotated using the box on the wall in the dome. The serial ports
were blown by lightning.
I’m trying to put together a fix.
- 29 OCT
It is OK to use the telescope again. Make sure you use the latest copy
of the checklist (in the notebook).
I replaced the declination motor. Pointing seems ok for now. Both RA and Dec drives need tuning, but I can schedule
that around labs.
- 16 OCT
TELESCOPE IS BROKEN!!!
DO NOT USE!!! I
broke it while trying to fix it.
The dec drive
will “run away” the minute you turn it on, so don’t turn it on!!
- 14 OCT
03 (true): Did our first multiple person training
run today at 4:30 with Jon Hakkila and Linwood
Wadsworth. Jon Hakkila wished to spread kudos to
those of us on making this telescope operation and control beautiful and
- 10 OCT
03 (neff): I will print out new copies of the
checklist before next week's training sessions.
- 10 OCT
03 (neff): I put START OBSERVING and STOP
OBSERVING on the start menu from the Observer account. I modified and
clarified the checklist. I put tape on the control box panel to indicate
the correct positions (i.e. away from the black tape) of each switch. Only
the Track swith and the Halt Motors button are
normally used, but the positions of the other switches are important. I
also put a carbon copy of the local documentation page on public site.
- 9 OCT
03 (neff): There is a new "Start
Observing" and "Stop Observing" icons on the desktop. When
called for in the checklist, double click on the "Start
Observing" icon. To "run" it, click the icon in the middle with
a vertical, downpointing arrow on it.
Alternatively, Macro->Run. When you are ready to quit, minimize (do not
exit) TheSky and double click on the "Stop
Observing" icon. Always answer "no" if it prompts you to
save startup.orc, shutdown.orc,
- 8 OCT
03 (neff): Updates to the Procedures Document. Updates to the links from
the main documentation page.
- 6 OCT
03 (neff): The first training session is
scheduled for Tuesday night, 7 October. Everyone will receive a copy of
the checklist and a copy of the policies. Make sure you check for updates
to these documents each time you observe.